OpenStudy (anonymous):

Factor the polynomial f(x)=x^3+4x^2+x-6 completely given that -3 is a zero of the polynomial ?

OpenStudy (anonymous):

x=-3 is zero so x+3 is a factor for polynomial\[f(x)=x^3+3x^2+x^2+3x-2x-6\]

OpenStudy (anonymous):

factor by grouping

OpenStudy (anonymous):

grouping?

OpenStudy (anonymous):

\[f(x)=x^3+3x^2+x^2+3x-2x-6\] -------- -------- -------- 3 group each with factor x+3

OpenStudy (anonymous):

for example\[x^2+3x=x(x+3)\]